Algebra:how much water must be boiled away to increase the concentration of 300 gallons of salt solution from 2% to 3%?
Answer by checkley77(12844)

.03*300-x=.02(300-x)
9-x=6-.02x
-x+.02x=6-9
-.98x=-3
x=-3/-.98
x=3.06 gal.
Proof:
.03*300-3.06=.02(300-3.06)
9-3.06=6-.02*3,06
5.94=6-.06
5.94=5.94